In a particular rural area of Victoria, 73% of learner drivers pass their driving test on the first attempt. If 150 learner drivers are selected, what is the probability that more than 120 drivers pass?
Answer by Boreal(15235)

This is a one-sample proportion test
What is the likelihood if the true proportion is 73%, that 80 or more per cent pass?
z=(0.80-0.73)/sqrt{(.73)(.27)/150}
z=0.07/0.0362=1.931
P=0.0267